Barry steroids company has $1000 par value bonds outstanding at 14% interest. The bonds will mature in 40 years. If the percent yield to maturity is 10%, what percent of the total bond value does the repayment of principle represent? Do not round intermediate calculations. And put your answer as a percent rounded to two decimal places.

Principal as a percentage of bond price =
Answer #1

bond price

=(1000*14%)*((1-(1+10%)^(-40))/10%)+1000/(1+10%)^40

=1391.16

principal amount present value=1000/(1+10%)^40=22.09

Principal as a percentage of bond price=22.09/1391.16=1.59%

the above is answer..